OpenStudy (anonymous):

Express the volume of a cone, V, as a function of its radius, r, if the radius is 1/11 of the height. V(r) =

OpenStudy (anonymous):

Volume of a cone is\[\[V(r,h)=\frac{1}{3} \pi r^3 h\]\]If the r= h/11, then\[V(r)=\frac{11}{3} \pi r^4\]

OpenStudy (anonymous):

what?